OpenStudy (brightemmerald):

What value in place of the question mark makes the polynomial below a perfect square trinomial? A. 22 B. 44 C. 121 D. 11

zepdrix (zepdrix):

The polynomial below?

OpenStudy (brightemmerald):

Sorry, x^2 +22x =?

zepdrix (zepdrix):

\[\large\rm x^2+bx\]Take half of your b-coefficient and square it,\[\large\rm \left(\frac{b}{2}\right)^2\]So for your polynomial, that means do those things to the 22, ya?

OpenStudy (brightemmerald):

So, (22/2)^2

zepdrix (zepdrix):

Yes.

OpenStudy (brightemmerald):

So, is that my final answer?

zepdrix (zepdrix):

Well that's not one of the answer choices :) lol So no. You'll have to simplify it. Cut 22 in half (that's 11, right?) then square it.

OpenStudy (brightemmerald):

121

zepdrix (zepdrix):

yayyyy good job \c:/

OpenStudy (brightemmerald):

(>0.0)>(^0.0^)<(0.0<)

zepdrix (zepdrix):

hehe :3